If the own-price elasticity of chilli supply is 0.40 and the chilli price is $2.20/lb., what would be the price of chilli if the quantity of cheese supplied increased from 1,250,000 lb to 1,275,000 lb?
Answer #1

Price elasticity of Supply = % change in Quantity supplied / Percentage change in Price

Now,

Original Quantity supplied: 1250,000 lb

New quantity supplied: 1275,000 lb

Change in Quantity supplied = 1275,000 -1250,000 = 25,000 lb

Percentage change in quantity supplied = Change in quantity supplied/ Original quantity supplied

                                                                = 25,000 lb /1250,000 lb = 2%

Now, Price elasticity of supply = 0.4

0.4 = Percentage change in quantity supplied i.e. 2% / percentage change in price

Therefore, Percentage change in price = 5% (price is increased as supply has also increased)

Original price: $2.20 /lb

Percentage increase in price = 5%

New price = $2.20 / lb + 5% = $ 2.31 per lb